OpenStudy (anonymous):

Given 10.8 grams of substance Y, if the substance absorbs 231 kilojoules of energy, and the specific heat of the substance is 6.55 J/g degrees Celsius, what is the change in temperature?

OpenStudy (chmvijay):

Q=M*C*Delta T u know Q=231Kja nd M=10.8 g ,C= specific heat of that material=6.55J/g find Delta T

OpenStudy (anonymous):

Is the answer 3.625? If so, how do you get there?

OpenStudy (chmvijay):

3.265 i think

OpenStudy (chmvijay):

Delta T=Q/M*C =231/10.8*6.55

OpenStudy (anonymous):

Can you show me the steps to get this? I get a much higher number because I converted 231 kJ to 231000 J. Is that not necessary?

OpenStudy (chmvijay):

yaaa u can convert it !!

OpenStudy (anonymous):

So, would the answer be 3,625 degrees Celsius? That seems awfully high

OpenStudy (chmvijay):

yaaa

OpenStudy (anonymous):

So which is right ?

OpenStudy (chmvijay):

to convert it into Joules is correct whatever the value may be

OpenStudy (chmvijay):

bcoz J/g*J/g here all units should cancel right

OpenStudy (anonymous):

The options I'm given are: a. -3.625 degrees Celsius b. 3.625 degrees Celsius c. 140 degrees Celsius d. -140 degrees Celsius I suppose that 3.625 is the closest to the right answer, right?

OpenStudy (chmvijay):

yaa its looks like that

OpenStudy (anonymous):

Okay, well thank you for all your help and your time in helping me! I appreciate it very much.

OpenStudy (chmvijay):

ur welcome you with such a good problems :)